# Keep it Light! Simplifying Image Clustering via Text-Free Adapters

### v1

@@ Abstract
Many competitive clustering pipelines have a multi-modal design, leveraging large language models (LLMs) or other text encoders, and text-image pairs, which are often unavailable in real-world downstream applications. Additionally, such frameworks are generally complicated to train and require substantial computational resources, making widespread adoption challenging. In this work, we show that in deep clustering, competitive performance with more complex state-of-the-art methods can be achieved using a text-free and highly simplified training pipeline. In particular, our approach, Simple Clustering via Pre-trained models (SCP), trains only a small cluster head while leveraging pre-trained vision model feature representations and positive data pairs. Experiments on benchmark datasets—including CIFAR-10, CIFAR-20, CIFAR-100, STL-10, ImageNet-10, and ImageNet-Dogs—demonstrate that SCP achieves highly competitive performance. Furthermore, we provide a theoretical result explaining why, at least under ideal conditions, additional text-based embeddings may not be necessary to achieve strong clustering performance in vision.
@@ 1 Introduction
Powered by the expressive capabilities of neural networks, deep clustering (DC) has redefined the state of the art (SOTA) in image clustering, outperforming traditional algorithms by indisputable margins. These DC pipelines uncover subtle features that can significantly enhance the power of downstream learners. A highly non-exhaustive list of algorithms redefining their respective SOTA includes DC-powered methods for community detection [1, 2], anomaly detection [3], image segmentation [4, 5], object detection [6, 7], and medical applications [8]. However, many recent DC pipelines rely on massive models, often powered by large language models (LLMs), which may be unnecessarily compute-intensive for clustering tasks, as well as complicated training schemes. This motivates our main research question:
Is SOTA performance achievable with a simple DC pipeline?
Indeed, we show that competitive performance, SOTA and near-SOTA, can be achieved using only a small adapter modifying the features generated by a pre-trained text-free, image encoder, i.e. not requiring any additional input text from the user, LLMs, or any other variants of text encoders. The result is a lightweight and simple end-to-end DC pipeline, which is computationally cheap enough to be run on a standard L4 GPU. Hence, our framework is easy to deploy in real-world tasks, as it is computationally accessible to most practitioners and does not require text-image pairs, which may not always be readily available in practical clustering applications.
@@ 1.1 Theoretical Motivation
The intuition behind why a text-free classifier should, at least theoretically, be able to match the power of a classifier using text embeddings together with the raw pixel data is rooted in our theoretical results. If $X$ represents a random image and $Z$ is its corresponding semantic description, which itself is a compressed representation of the information only in the random image $X$, then any classification task $Y$ that depends on both the image $X$ and its textual information can ultimately be expressed as a function of the image $X$ alone. We denote the $\sigma$-algebra generated by the random variable $X$ by $\sigma(X)$.
Proposition 1.1 (Lossless Amortization Principle (LAR)).
@@
Proof.
See Appendix B. ∎
Further, the proposition above seems to align with the Platonic Representation Hypothesis proposed in [9], which suggests that latent representations in deep networks converge despite being trained on distinct data modalities. This supports our claim that, as long as the backbone vision model provides sufficiently good latent representations, the addition of extra modality inputs may be unnecessary.
Note that we acknowledge that, generally speaking, using multimodal representations and additional information for our networks to make predictions empirically tends to yield better results. Rather than disputing this, we aim to theorize whether, in principle and under ideal conditions, clustering could be achieved using the simplest model possible, without cumbersome multimodal data collection.
Our LAR principle (Proposition 1.1) shows that any semantically-powered image classifier may be realized by a text-free classifier “$F$”. However, can this theoretical classifier be practically realized, at-least approximately? Our main theoretical result guarantees that the theoretical text-free classifiers can indeed be approximately implemented by an MLP with a $\operatorname{softmax}$ output activation. Our guarantee holds for multiclass MLP classifiers using any standard activation function such as $\operatorname{ReLU}$, $\operatorname{Swish}$, or $\operatorname{softplus}$.
Theorem 1.2 (Text-Free DC is Powerful Enough).
In the setting of Proposition 1.1, let $\rho:\mathbb{R}\to\mathbb{R}$ be an activation function with at least one point of continuous and non-zero differentiability. Then, for every $0<\varepsilon\leq 1$, there is an MLP $\hat{F}:\mathbb{R}^{d}\to\mathbb{R}^{C}$, with activation function $\rho$, satisfying
@@
Proof.
See Appendix B. ∎
@@ 1.2 Simplifying Contemporary Deep Clustering
Modern clustering pipelines, e.g. [10, 11] usually involve: (i) learning powerful initial latent representations through self-supervised techniques, e.g. joint-embedding approaches [12, 13], and (ii) gradually refining the representation and clustering membership by minimizing an objective function [14].
@@
The emergence of large-scale pre-trained models optimized according to a contrastive learning objective using both natural language such as CLIP [15] and purely image-based encoders like DINO [16, 17], has proven highly effective in capturing rich, general-purpose feature representations for a variety of downstream tasks. Most SOTA-achieving DC pipelines leverage these powerful representations alongside text-embedding [18, 19] (the $Z$ in Theorem 1.2) for effective latent feature alignment. However, as we will see in Section 4, these text-based representations seem to contain redundant features, given that similar performance can be achieved without them. As we have already mentioned, being able to eliminate the need for text would make downstream applications, particularly in data-scarce regimes, much more accessible to the end user.
@@
Typically, modern DC pipelines are trained using self-supervised methods by optimizing contrastive losses [20]. While contrastive methods rely on both positive and negative pairs, other approaches like BYOL [21] have demonstrated that it is possible to learn effective representations without negative pairs by simply maximizing the agreement between two views of the same data. This suggests that maximizing similarities of positive pairs alone can yield high-quality representations. We will follow a similar approach in this work too, see Section 3.
@@
Even if the model is approximately optimal and its parameters are perfectly adjusted, it is not clear that any real-world unsupervised training procedure can achieve the theoretical expressivity guaranteed by Theorem 1.2. Hence, we would like to empirically answer whether unimodal text-free DC pipelines can rival multimodal ones. We would like to emphasize that obtaining powerful DC performance, with the additional requirement that our DC pipeline is simple, unimodal, and text-free, is highly nontrivial since even the current SOTA in deep clustering exhibits a substantial performance gap with respect to supervised image classification models. Indeed, supervised methods typically achieve accuracies well-above $90\%$ on image datasets, while unsupervised clustering approaches rarely surpassed 50% clustering accuracy on CIFAR-20, and many fall below $50\%$ [22, 23, 24, 25]. This is because most unsupervised clustering pipelines struggle with complex natural images due to high intra-class variability.
@@
To address the challenges discussed so far, we propose Simple Clustering via Pre-trained models (SCP). Our main contributions are as follows:
  • Simple DC Pipeline: Our method yields competitive results while maintaining a simple yet effective architecture, requiring no additional feature layers, support sets, teacher-student networks, text information, or exponential moving averages.
  • Text-Free (Unimodal) Encodings: We integrate the powerful image encoder from CLIP (and also experiment with DINO) into a self-supervised clustering framework and show that this approach achieves competitive clustering performance on benchmark datasets such as CIFAR-10 [26], CIFAR-20 [26] and STL-10 [26] to current SOTA.
  • Principled Approach: Our text-free approach is principled, based on the approximate representation result previously presented in Theorem 1.2, which guarantees that text-free embeddings are theoretically enough.
@@ 2.1 Self-Supervised Learning
Self-supervised learning learns representations from data without explicit labels. The objective is to create a representation space where positive pairs are closer together, while negative pairs are pushed farther apart [27].
SimCLR [12] uses data augmentations, such as flipping and colour jittering, to create positive and negative pairs for optimizing objectives. It also introduces a projection head that maps embeddings into a space where contrastive loss is applied. BYOL [21] shows that high-quality representations can be learned by simply maximizing agreement between two augmented views of the same input, without requiring negative pairs. Building on these advancements, SimSiam [28] eliminates the need for both negative pairs and momentum encoders by introducing a stop-gradient operation, which effectively prevents representational collapse. Inspired by these methods, we adopt similar ideas to develop a simple and effective self-supervised framework for image clustering.
@@ 2.2 Pre-trained Models in Vision
Building on advances in self-supervised learning, CLIP [15] introduced a paradigm of contrastive pre-training that aligns images with corresponding textual descriptions. This approach enables broad task generalization without task-specific fine-tuning. DINO [16], which stands for self-distillation with no labels, demonstrates a self-supervised method for optimizing a student network from a teacher network based on vision input data only.
One of the key advantages of pre-trained models like CLIP is their ability to eliminate the need for training models from scratch for downstream tasks, significantly reducing computational costs and time. Instead of training a self-supervised neural network from the ground up, pre-trained models provide high-quality feature representations out of the box, leading to faster experimentation and improved performance on a variety of tasks. The scalability of CLIP has been further validated by openCLIP [29], which extended CLIP using the larger Vision Transformer models [30]. Similarly, models such as DINO [31] and DINOv2 [17] are capable of processing visual data and mapping it to high-quality latent representations.
@@ 2.3 Image Clustering via Pre-trained Models
To address the challenges of scaling to modern image datasets, methods such as NMCE [32] and MLC [33] have integrated deep learning with manifold clustering using the minimum coding rate principle [34]. Building on this idea, CPP [35] further refines CLIP features and estimates the optimal number of clusters when unknown. TEMI [36] improves clustering by leveraging associations between image features, introducing a variant of pointwise mutual information with instance weighting. Unlike our approach, TEMI utilizes a nearest-neighbors set and an exponential moving average for parameter optimization.
SIC [19] leverages multi-modality by mapping images to a semantic space and generating pseudo-labels based on image-semantic relationships. More recently, TAC [18] utilizes the textual semantics of WordNet [37] to enhance image clustering by selecting and retrieving nouns that best distinguish the images, facilitating collaboration between text and image modalities through mutual cross-modal neighborhood distillation.
Current pre-trained approaches often rely on heavy or complex architectures to ensure consistency, motivating us to develop a simple yet effective pipeline for image clustering. Our method requires only a simple clustering head and basic data augmentations, demonstrating strong competitiveness among recent models.
@@ 3 Our Method
Recent methods such as CC [23] and CPP [35] decouple the latent space into clustering and feature spaces. TAC [18] utilizes text information, and TEMI [36] employs self-distillation networks to enhance clustering. In contrast, our approach remains simple and efficient without relying on these techniques. As shown in Fig. 1, SCP consists of only two components: a pre-trained frozen backbone for pair construction, denoted as $f(.)$, and a trainable cluster head $g(.)$.

### v2

@@ Abstract
In the era of pre-trained models, effective classification can often be achieved using simple linear probing or lightweight readout layers. In contrast, many competitive clustering pipelines have a multi-modal design, leveraging large language models (LLMs) or other text encoders, and text-image pairs, which are often unavailable in real-world downstream applications. Additionally, such frameworks are generally complicated to train and require substantial computational resources, making widespread adoption challenging. In this work, we show that in deep clustering, competitive performance with more complex state-of-the-art methods can be achieved using a text-free and highly simplified training pipeline. In particular, our approach, Simple Clustering via Pre-trained models (SCP), trains only a small cluster head while leveraging pre-trained vision model feature representations and positive data pairs. Experiments on benchmark datasets, including CIFAR-10, CIFAR-20, CIFAR-100, STL-10, ImageNet-10, and ImageNet-Dogs, demonstrate that SCP achieves highly competitive performance. Furthermore, we provide a theoretical result explaining why, at least under ideal conditions, additional text-based embeddings may not be necessary to achieve strong clustering performance in vision.
Index Terms:
Image Clustering; Deep Clustering; Self-supervised Learning.
@@ 1 Introduction
Powered by the expressive capabilities of neural networks, deep clustering (DC) has redefined the state of the art (SOTA) in image clustering, outperforming classical algorithms by indisputable margins. Modern DC pipelines uncover subtle feature structures that substantially enhance the effectiveness of downstream learners. A highly non-exhaustive list of domains where DC-powered algorithms have reshaped the SOTA includes community detection [saeedi2018novel, yaroslavtsev2018massively], anomaly detection [pang2021deep], image segmentation [minaee2021image, liang2023clustsegclusteringuniversalsegmentation], object detection [zhao2019object, kim2024garfieldgroupradiancefields], and various medical applications [zhao2020joint].
However, in addition to image encoders, many recent DC pipelines rely on massive models—often involving large language models (LLMs)—that are unnecessarily compute-intensive for clustering tasks and introduce complicated training schemes. Unlike supervised learning methods, which typically require only a lightweight readout layer or linear probing for classification, current unimodal deep clustering frameworks frequently incorporate additional components such as support sets, feature projection layers, or teacher–student architectures. While these designs have achieved impressive performance, they also introduce extra hyperparameters, increase training complexity, and ultimately raise an important research question.
1.1 Is SOTA performance achievable with a simple DC pipeline?
As image clustering often serves as a foundational step for a variety of downstream tasks, a simple yet effective solution can thus provide broad utility to the community. Indeed, we show that competitive performance, SOTA and near-SOTA, can be achieved using only a small adapter modifying the features generated by a pre-trained text-free, image encoder, i.e. not requiring any additional input text from the user, LLMs, or any other variants of text encoders. The result is a lightweight and simple end-to-end DC pipeline, which is computationally cheap enough to be run on a standard L4 GPU. In parallel to how linear probing or read-out layers are deployed in classification tasks, we show that our adapter is easy to adopt in real-world applications, as it is simple and effective.
@@ 1.2 Theoretical Motivation
Text-image pairs may not always be readily available in practical clustering applications. The intuition behind why a text-free classifier should, at least theoretically, be able to match the power of a classifier using text embeddings together with the raw pixel data is rooted in our theoretical results. If $X$ represents a random image and $Z$ is its corresponding semantic description, which itself is a compressed representation of the information only in the random image $X$, then any classification task $Y$ that depends on both the image $X$ and its textual information can ultimately be expressed as a function of the image $X$ alone. We denote the $\sigma$-algebra generated by the random variable $X$ by $\sigma(X)$.
Proposition 1 (Lossless Amortization Principle (LAR)).
@@
Proof of Proposition 1.
By the Doob–Dyknin lemma, see e.g. [KallenbergProbability_2021, Lemma 1.14], since $Z$ is $\sigma(X)$-measurable iff there exists some Borel measurable function $g:(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))\to(\mathbb{R}^{D},\mathcal{B}(\mathbb{R}^{D}))$ such that
$Z=g(X).$ (2)
Define $F:(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))\to(\mathbb{R}^{n},\mathcal{B}(\mathbb{R}^{n}))$ by composition:
$F(x)\triangleq f(x,g(x)).$ (3)
Then, by (2) and (1),
$Y=f(X,Z)=f(X,g(X))=F(X),$
which establishes the claim. ∎
Further, the proposition above seems to align with the Platonic Representation Hypothesis proposed in [pmlr-v235-huh24a], which suggests that latent representations in deep networks converge despite being trained on distinct data modalities. This supports our claim that, as long as the backbone vision model provides sufficiently good latent representations, strong performance can be achieved through a simple DC pipeline without relying on additional modality inputs.
We acknowledge that, generally speaking, incorporating multimodal representations and additional information in our network tends to yield better empirical performance. Rather than disputing this, we deliberately refrain from using such frameworks, and instead aim to explore whether, in principle and under ideal conditions, clustering could be achieved using the simplest model possible, without multimodal data collection.
Our LAR principle (Proposition 1) shows that any semantically-powered image classifier may be realized by a text-free classifier “FF”. However, can this theoretical classifier be practically realized, at-least approximately? Our main theoretical result guarantees that the theoretical text-free classifiers can indeed be approximately implemented by an MLP with a $\operatorname{softmax}$ output activation. Our guarantee holds for multiclass MLP classifiers using any standard activation function such as $\operatorname{ReLU}$, $\operatorname{Swish}$, or $\operatorname{softplus}$.
Theorem 1 (Text-Free DC is Powerful Enough).
In the setting of Proposition 1, let $\rho:\mathbb{R}\to\mathbb{R}$ be an activation function with at least one point of continuous and non-zero differentiability. Then, for every $0<\varepsilon\leq 1$, there is an MLP $\hat{F}:\mathbb{R}^{d}\to\mathbb{R}^{C}$, with activation function $\rho$, satisfying
@@
Proof of Theorem 1.
Let $\mu\triangleq X_{\#}\mathbb{P}$ denote the law of $X$. By [KlenkeBook_2020, Theorem 13.6], $\mu$ is a Radon measure. Since $\mathbb{R}^{d}$ is second-countable and locally compact, Lusin’s theorem [KlenkeBook_2020, Exercise 13.1.3] yields, for every $\varepsilon\in(0,1]$, a compact set $\mathcal{K}_{\varepsilon}\subset\mathbb{R}^{d}$ such that
$\mu(\mathcal{K}_{\varepsilon})\geq 1-\varepsilon,$ (5)
and $F$ is continuous on $\mathcal{K}_{\varepsilon}$.
Now, let $\rho$ satisfy [kratsios2022universal, Assumption 1] (as in [kidger2020universal]), and recall that the softmax activation satisfies [kratsios2022universal, Assumption 8]. Then, the universal approximation theorem [kratsios2022universal, Theorem 37(ii)] ensures the existence of an MLP $\hat{F}:\mathbb{R}^{d}\to\mathbb{R}^{C}$ such that
$\sup_{x\in\mathcal{K}_{\varepsilon}}\big|F(x)-\operatorname{softmax}\circ\hat{F}(x)\big|<\varepsilon.$ (6)
Define
$B\triangleq\bigl\{x\in\mathbb{R}^{d}:\,|F(x)-\operatorname{softmax}\circ\hat{F}(x)|\geq\varepsilon\bigr\}.$
Then $\mathcal{K}_{\varepsilon}\subseteq\mathbb{R}^{d}\setminus B$. Let $L(x)=|F(x)-\operatorname{softmax}\circ\hat{F}(x)|$. Since $L$ is measurable, so is $B=L^{-1}([\varepsilon,\infty))$.
Combining (5) and (6), we get
$\begin{aligned} \mathbb{P}\!\Big(|F(X)-\operatorname{softmax}\!\circ\!\hat{F}(X)|<\varepsilon\Big)&=\mu\!\Big(|F(x)-\operatorname{softmax}\!\circ\!\hat{F}(x)|<\varepsilon\Big)\\ &=\mu(\mathbb{R}^{d}\setminus B)\\ &\geq\mu(\mathcal{K}_{\varepsilon})\\ &\geq 1-\varepsilon.\end{aligned}$
Finally, by Proposition 1 and (1),
$F(X)=f(X,Z).$ (7)
Substituting (7) into (1.2) gives
$\mathbb{P}\!\Big(|f(X,Z)-\operatorname{softmax}\!\circ\!\hat{F}(X)|<\varepsilon\Big)$ $\geq 1-\varepsilon.$ (8)
which concludes the proof. ∎
@@ 1.3 Simplifying Contemporary Deep Clustering
Modern clustering pipelines, e.g. Van et al. and Ding et al. [van2020scan, ding2023unsupervised] usually involve: (i) learning powerful initial latent representations through self-supervised techniques, e.g. joint-embedding approaches [chen2020simple, he2020momentum], and (ii) gradually refining the representation and clustering membership by minimizing an objective function [caron2019deepclusteringunsupervisedlearning].
@@
The emergence of large-scale pre-trained models optimized according to a contrastive learning objective using both natural language such as CLIP [radford2021learning] and purely image-based encoders like DINO [caron2021emerging, oquab2024dinov2learningrobustvisual], has proven highly effective in capturing rich, general-purpose feature representations for a variety of downstream tasks. Most SOTA-achieving DC pipelines leverage these powerful representations alongside text-embedding [li2023image, cai2023semantic] (the $Z$ in Theorem 1) for effective latent feature alignment. However, it would involve extra computations and we will see in Section 4, these text-based representations seem to contain redundant features, given that comparable performance can be achieved with a unimodal framework. As we have already mentioned, being able to eliminate the need for text would make downstream applications, particularly in data-scarce regimes, much more accessible to the end user.
@@
Typically, modern DC pipelines are trained using self-supervised methods by optimizing contrastive losses [oord2018representation]. While contrastive methods rely on both positive and negative pairs, other approaches like BYOL [grill2020bootstrap] have demonstrated that it is possible to learn effective representations without negative pairs by simply maximizing the agreement between two views of the same data. This suggests that maximizing similarities of positive pairs alone can yield high-quality representations. We will follow a similar approach in this work too, see Section 3.
@@
Even if the model is approximately optimal and its parameters are perfectly adjusted, it is not clear that any real-world unsupervised training procedure can achieve the theoretical expressivity guaranteed by Theorem 1. Hence, we would like to empirically answer whether a simple DC pipelines can rival both multimodal methods and its unimodal competitors. We would like to emphasize that obtaining powerful DC performance, with the additional requirement that our DC pipeline is simple, and text-free, is highly nontrivial since even the current SOTA in deep clustering exhibits a substantial performance gap with respect to supervised image classification models. Indeed, supervised methods typically achieve accuracies well-above $90\%$ on image datasets, while unsupervised clustering approaches rarely surpassed 50% clustering accuracy on CIFAR-20, and many fall below $50\%$ [dang2021nearest, li2021contrastive, shen2021you, niu2022spice]. This is because most unsupervised clustering pipelines struggle with complex natural images due to high intra-class variability.
@@
To address the challenges discussed so far, we propose Simple Clustering via Pre-trained models (SCP), particularly in scenarios where text information or complex frameworks are not accessible. Much like linear probing in classification, SCP provides a simple yet powerful way to leverage pre-trained visual features for accessible deep clustering. Our main contributions are as follows:
  • Simplicity: Our method yields competitive results while maintaining a simple yet effective architecture, requiring no additional feature layers, support sets, teacher-student networks, text information, or exponential moving averages during training.
  • Applicability: Our method seamlessly integrates the powerful image encoder from various pre-trained models such as CLIP and DINO into a self-supervised clustering framework. We show that this text-free adapter achieves highly competitive clustering performance on benchmark datasets such as CIFAR-10 [krizhevsky2009learning], CIFAR-20 [krizhevsky2009learning] and STL-10 [krizhevsky2009learning] to current SOTA.
  • Principled Approach: Our text-free approach is principled, based on the approximate representation result previously presented in Theorem 1, which guarantees that text-free embeddings are theoretically enough.
@@ 2.1 Self-Supervised Learning
Self-supervised learning learns representations from data without explicit labels. The objective is to create a representation space where positive pairs are closer together, while negative pairs are pushed farther apart [geiping2023cookbook].
SimCLR [chen2020simple] uses data augmentations, such as flipping and colour jittering, to create positive and negative pairs for optimizing objectives. It also introduces a projection head that maps embeddings into a space where contrastive loss is applied. BYOL [grill2020bootstrap] shows that high-quality representations can be learned by simply maximizing agreement between two augmented views of the same input, without requiring negative pairs. Building on these advancements, SimSiam [chen2020exploringsimplesiameserepresentation] eliminates the need for both negative pairs and momentum encoders by introducing a stop-gradient operation, which effectively prevents representational collapse. Inspired by these methods, we adopt similar ideas to develop a simple and effective self-supervised framework for image clustering.
@@ 2.2 Pre-trained Models in Vision
Building on advances in self-supervised learning, CLIP [radford2021learning] introduced a paradigm of contrastive pre-training that aligns images with corresponding textual descriptions. This approach enables broad task generalization without task-specific fine-tuning. DINO [caron2021emerging], which stands for self-distillation with no labels, demonstrates a self-supervised method for optimizing a student network from a teacher network based on vision input data only. The scalability of CLIP has been further validated by openCLIP [Cherti_2023], which extended CLIP using the larger Vision Transformer models [dosovitskiy2020image]. Similarly, models such as DINO [9709990] and DINOv2 [oquab2024dinov2learningrobustvisual] are capable of processing visual data and mapping it to high-quality latent representations.
@@ 2.3 Image Clustering via Pre-trained Models
To address the challenges of scaling to modern image datasets, methods such as NMCE [li2022neural] and MLC [deng2023acp] have integrated deep learning with manifold clustering using the minimum coding rate principle [Arthur_Vassilvitskii_2007]. Building on this idea, CPP [chu2024image] further refines CLIP features and estimates the optimal number of clusters when unknown. TEMI [adaloglou2023exploring] improves clustering by leveraging associations between image features, introducing a variant of pointwise mutual information with instance weighting. Unlike our approach, TEMI utilizes a nearest-neighbors set and an exponential moving average for parameter optimization.
SIC [cai2023semantic] leverages multi-modality by mapping images to a semantic space and generating pseudo-labels based on image-semantic relationships. More recently, TAC [li2023image] utilizes the textual semantics of WordNet [miller1995wordnet] to enhance image clustering by selecting and retrieving nouns that best distinguish the images, facilitating collaboration between text and image modalities through mutual cross-modal neighborhood distillation.
Current pre-trained approaches often rely on text or complex architectures to ensure consistency, motivating us to develop a simple yet effective pipeline for image clustering. Our method requires only a lightweight clustering head and basic data augmentations, demonstrating competitiveness, simplicity, and applicability among recent models.
@@ 3 Our Method
Recent methods such as CC [li2021contrastive] and CPP [chu2024image] decouple the latent space into clustering and feature spaces. SIC [cai2023semantic] and TAC [li2023image] utilize text information, and TEMI [adaloglou2023exploring] employs self-distillation networks to enhance clustering. Despite the promising results, these methods involve additional components and complicated training as suggested by the table I below.
TABLE I: Summary of tunable hyperparameters and their concise roles in recent deep clustering methods. Fixed hyperparameters are omitted.
Methods Hyperparameters
CPP [chu2024image] $K_{\max}$, $\gamma$, $\varepsilon$
max # clusters; entropy etc.
TEMI [adaloglou2023exploring] $C$, $\beta$, $k$, $m$
# clusters; PMI trade-off etc.
TAC [li2023image] $K$, $\tilde{N}$, $\hat{\tau}$, $\hat{N}$, $\alpha$
# clusters; center size etc.
SIC [cai2023semantic] $c$, $\gamma_{u}$, $\gamma_{r}$, $\xi_{c}$, $\xi_{a}$, $k$, $\lambda$, $\beta$, $\tau_{m}$
# clusters; noun pruning etc.
SCP (Ours) $K$, $\alpha$
# clusters; loss weight
In contrast, in supervised learning, a lot of work perform classification only need a read-out layer or a simple linear-probing. Our approach try to keep the similar structure in the clustering as well and remains simple and efficient without relying on extra complexity as much as possible. As shown in Fig. 1, SCP involves only two tunable hyperparameters and comprises two components: a pre-trained frozen backbone for pair construction, denoted as $f(.)$, and a trainable cluster head $g(.)$.